# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v5

@@ 4.1.2 The Uncertainty Resolution Paradox and its Resolution
When LLM combines knowledge elements to fabricate facts or produce emergent insights in a given fixed context, it seems that
$\mu_{C}(A\cup B|Q)>\mu_{C}(A|Q)+\mu_{C}(B|Q),$
so novel insights may come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume subadditivity?
First, subadditivity of semantic information measure establishes a baseline where uncontrolled synergy of knowledge cannot appear. We remove from our analytical toolkit the potential measurement bias of information surplus that any information aggregation mechanism might generate.
Second, we prove that hallucinations are inevitable even when knowledge is subadditive (within a fixed context). In that setting, by Theorem 4, reasoning does not create new information. It is a transformation that reveals information already present in any knowledge representation. Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination. However, we should take care of a paradox those claims apparently bring into existence.
@@
But, if reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable. Doesn’t it mean the reasoning must have created new information, contrary to what we have just claimed? It seems we have a paradox.
To resolve it, we need the emergence operator $\mathcal{E}_{C}(X|Q)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
Notice that for all practical purposes the information measurement $\mu_{C}(X|Q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time) in a given context. Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X|Q)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X|Q)\leq\mu_{C}(\mathcal{E}_{C}(X|Q)|Q)$. That is possible only with organized information providing knowledge for reasoning. These are precisely the statements of Theorem 1 and Theorem 4.
@@
All there is to know is immediately accessible when we have unlimited computing resources (or time).
@@
$\theta(0)$ $=[1,1,1,0,0]\text{ (initial knowledge state)},$
$\theta(1)$ $=\mathcal{E}_{C}(\theta(0)|Q)=[1,1,1,1,1]\text{ (state after reasoning)},$
$y(0)$ $=\mu_{C}(\theta(0)|Q)=3\text{ bits},$
$y(1)$ $=\mu_{C}(\theta(1)|Q)=5\text{ bits}.$
@@
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)|Q)|Q)=\mu_{\infty}(\theta(0)|Q)$
and no new information is created.
In the next section, we relate that important property of response generation, namely, how much additional semantic information each LLM inference path can contribute, with the Semantic Information Conservation Principle.
4.1.3 Knowledge Representation in LLMs
The above Knowledge Representation finds its direct counterpart in the structure and state of a transformer architecture of LLMs.
The abstract knowledge space $\mathcal{K}$ is the high-dimensional embedding space of the model. Every fact, concept, and linguistic pattern the model has learned is represented by a specific element of that space. The metric $d_{\mathcal{K}}$ is typically implemented as cosine similarity or Euclidean distance, where proximity between vectors signifies a closer semantic relationship. Model parameters $\mathcal{W}$ correspond to the full set of trained weights in the transformer, including the attention matrices ($W_{Q},W_{K},W_{V},W_{O}$) and the matrices of the MLP layers. The mapping $\Phi(W)$ represents the totality of information, patterns, and factual associations stored within these trained weights. Activation patterns $\mathcal{A}$ correspond to the state of the residual stream during a single forward pass for a given input. An activation pattern is therefore a specific trajectory through the network’s computational graph. Finally, the mapping $\Psi(a_{i})$ represents the information made computationally accessible during a specific inference step and encoded in the residual stream.
The process of information processing under computational constraints maps to the sequential forward pass transformer inference.
The idea of computational budget $C$ can be represented by the depth or size of the network, e.g.,the number of transformer blocks. A deeper network has a larger budget $C$, allowing for more reasoning steps. In the context of multi-step reasoning, $C$ can also represent the number of iterations allowed in a Chain-of-Thought (CoT) sequence.
The semantic information measure $\mu_{C}$ can be linked to the model’s objective function. For a given input $A$, $\mu_{C}(A)$ can be understood as the measurable reduction in cross-entropy loss that the model achieves by processing $A$ with a budget $C$.
Finally, the emergence operator $\mathcal{E}_{C}$ can be identified with a transformer block (one multi-head attention layer followed by one MLP layer). This block applies a computational transformation $\mathcal{E}_{C}$ to the residual stream at layer $t$, producing the updated state at layer $t+1$, making information that was latent in the input explicitly accessible.
@@ 4.2 Statements, Queries, and Ground Truth
Queries and responses are statements expressed in natural language that LLM is able to process and generate. We need to compare those statements and relate them to knowledge. It is also useful to distinguish queries from responses by their nature, respectively, interrogative or declarative. Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions). We should also point out that we limit our study to the case of non-trivial queries that require integrating responses from many LLM inference paths. The following formal definition encapsulates those ideas.
Definition 8 (Statement Space).
The space of natural language statements $\mathcal{L}$ is a metric space with distance function $d_{\mathcal{S}}$. We distinguish query space $\mathcal{Q}\subset\mathcal{L}$ of interrogative statements and response space $\mathcal{R}\subset\mathcal{L}$ of declarative statements. Function $K:\mathcal{L}\rightarrow\mathcal{P}(\mathcal{K})$ maps statements to referenced knowledge. We call query $q$ non-trivial, if it is integrated from multiple knowledge subsets.
We must now rigorously define the ground truth mapping and establish formal relationships between the knowledge space $\mathcal{K}$ and response space $\mathcal{R}$.
For that, we should start by recognizing that the very idea of ground truth validation is limited (or subjective at least). It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query. Indeed, that limit is given by the Goedel’s Incompleteness Theorems for the formalized mathematical theories. Therefore, in practice that inaccessibility of the complete truth validation function (even in the idealized and rigorous mathematical reasoning) is one of the reasons why hallucination cannot be eliminated entirely. High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth. Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping, which we define as follows.
@@
The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$ associates each query $q\in\mathcal{Q}$ with knowledge elements constituting complete, correct answers:
[…]
In other words, $T(q)$ represents our best available approximation to (labeled as) true relevant knowledge. We accept the fact that there is some external agency telling us what is and what is not true answer to a query, and that evaluation has a form of label assigned to knowledge elements. But, for all that, we also focus on the queries that are non-trivial in the following sense.
@@
A query $q\in\mathcal{Q}$ is non-trivial with respect to computational budget $C$ if there exist knowledge configurations $\theta_{i}$ and $\theta_{j}$ for the context $Q=K(q)\cup T(q)$, such that
@@
A query is non-trivial when at least two agents disagree on some fact that is both encoded and available, and relevant to answering that query. For trivial queries impossibility results do not hold. That definition is important for the development of impossibility results. It measures the ability of a response generation system to store knowledge and expand knowledge that is necessary for the impossibility to emerge. We may view that ability to respond to non-trivial queries as an analog of the formal system complexity in the Goedel’s framework.
4.3 Auction-Theoretic Model of LLM Inference
There are many ways LLM can generate answers to a query. By design, it is a controlled random process governed by fine-tuned probability distributions. We are going to model that process as an auction in which agents (representing competing ideas is the space of all possible activation patterns $\mathcal{A}$) use their privately held knowledge (internal representations in $\mathcal{W}$) to construct bids (candidate responses) and submit them to the LLM response mechanism. That response mechanism then selects the best bid as a response to a query.
@@ 4.3 Auction-Theoretic Model of LLM Inference
  • Strategy space $S=S_{1}\times\cdots\times S_{n}$, where $S_{i}\subset\mathcal{P}(\mathcal{R})$,
  • Outcome function $g:S\rightarrow\mathcal{R}$ mapping strategies to responses,
  • Information contribution function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs of limiting hallucinations.
A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge $\theta_{i}\in\Theta_{i}\subset\mathcal{K}$ in response to a query. It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction. Examples of strategies include activations deciding which tokens to attend or how strongly to activate neurons for different patterns. Then, given a set of submitted bids (in the form of strategy profile) $s=(s_{1},\dots,s_{n})$, the response mechanism uses outcome function $g(s)$ to select and generate response. One straightforward example is a softmax function in LLM’s final layer. Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, information contribution function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output. In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
@@
The inference process involves $n$ agents $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$ competing by submitting bids to mechanism $\mathcal{M}$. Each agent $a_{i}$ constructs bid $s_{i}(\theta_{i})$ based on private knowledge $\theta_{i}\in\Theta_{i}$ to maximize utility:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta),$ (18)
where $v_{i}\geq 0$ is valuation and $p_{i}$ is information contribution. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
One way to think about the concept of agents is that it represents different attention heads in transformer models, or components of mixture-of-experts architectures, or activation paths fired by the same query. Each agent’s private knowledge represents model parameters utilized by an agent to construct a response. The utility function can be best described as a partial contribution of a candidate response to LLM loss function minimization. It is designed to explain why any inference path is activated in response to the LLM prompt. For example, negative log likelihood loss can be decomposed with respect to attention heads, and linearized to measure how much attention head contributes to token prediction.
Let us point out that agents represent algorithms and architectural components of any LLM implementation, they are not conscious beings with individual goals. That is important, as it refers to the Revelation Principle in game theory. If we know strategies (of using knowledge) the players (conscious beings with individual goals) will use in a game, we can implement those strategies as an algorithm (or agent) with parameters representing private preferences or knowledge of the players. In the case considered, we are the players selecting architecture and training the LLM.
We assume that individual utility depends on all responses $s=(s_{1},\dots,s_{n})$ and knowledge $\theta=(\theta_{1},\dots,\theta_{n})$ used by all agents, $u_{i}=u_{i}(s,\theta)$. Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and information contribution $p_{i}=p_{i}(s,\theta)$ value assigned.
The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $\nabla_{\theta_{i}}L_{W}$).
@@
We can also illustrate the concept of utility function with the following example:
$u_{i}(s,\theta)$ $=\underbrace{\alpha\cdot\text{Accuracy}(g(s),\theta_{i})}_{\text{Valuation}}-\underbrace{\gamma\cdot\text{Contradiction}(s_{i},\theta_{i})}_{\text{Consistency Cost}}.$
$\text{Accuracy}(g(s),\theta_{i})$ measures how well the response aligns with the knowledge $\theta_{i}$ agent $a_{i}$ is using. $\text{Contradiction}(s_{i},\theta_{i})$ measures how much the strategy $s_{i}$ conflicts with internal knowledge state $\theta_{i}$ (both extracted from $s$ and $\theta$ profiles).
Independent private knowledge assumption
One of the critical and controversial assumptions we make is is that knowledge profiles $\theta_{i}$ are independently distributed. It should be noted that in transformers, heads may share layer-norm signals, residual streams, and optimizer state. That may result in correlated (or, so called, affiliated) inference signals between token representations. Furthermore, the superposition hypothesis, introduced in mechanistic interpretability studies, suggests cooperative error-canceling due to strong correlations.
To address that issue, we show that impossibility arises in both independent and correlated settings in general and in the LLM transformer environment. Namely, we refer to Green-Laffont theorem green1977characterization in the independent private knowledge setting. Then, we refer to the truthful probability elicitation and proper scoring theory introduced by Savage savage1971elicitation , which does not require independent private knowledge.

### v7

@@ 4.1.5 The Uncertainty Resolution Paradox and its Resolution
As we have just seen, when inference combines knowledge elements to fabricate facts or produce emergent insights in a given fixed context, it seems that
$\mu_{C}(A\cup B|Q)>\mu_{C}(A|Q)+\mu_{C}(B|Q).$ (39)
Novel insights may come from existing data. That is the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume subadditivity of the semantic information measure instead?
First, subadditivity establishes a baseline where uncontrolled synergy of knowledge cannot appear. We remove from our analytical toolkit the potential bias that any information aggregation mechanism might generate.
Second, we prove that hallucinations are inevitable even when knowledge is subadditive (within a fixed context). It is in that setting that Theorem 5 states that reasoning does not create new information. It can only reveal information already present in any knowledge representation.
Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination. To explain why, we should take care of the following paradox.
@@
But, if reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable. Doesn’t that mean the reasoning must have created new information, contrary to what we have just claimed? It seems we have a paradox.
To resolve it, we need the emergence operator $\mathcal{E}_{C}$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
Notice that for all practical purposes the information measurement $\mu_{C}(X|Q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time) in a given context. Therefore, given the result of computationally bounded reasoning, we can observe information gain through the computational work, i.e., $\mu_{C}(X|Q)\leq\mu_{C}(\mathcal{E}_{C}(X|Q)|Q)$. That is possible only with organized information providing knowledge for reasoning.
@@
By Theorem 3, all there is to know is immediately accessible when we have unlimited computing resources (or time).
Example.
@@
$\theta(0)$ $=[1,1,1,0,0]\text{ (initial knowledge state)},$ (42)
$\theta(1)$ $=\mathcal{E}_{C}(\theta(0)|Q)=[1,1,1,1,1]\text{ (state after reasoning)},$ (43)
$y(0)$ $=\mu_{C}(\theta(0)|Q)=3\text{ bits},$ (44)
$y(1)$ $=\mu_{C}(\theta(1)|Q)=5\text{ bits}.$ (45)
@@
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)|Q)|Q)=\mu_{\infty}(\theta(0)|Q)$ (46)
and no new information is created. ∎
Axiomatic Reasoning
In axiomatic reasoning we can admit reasoning rules $f\in\mathcal{F}_{C}$ such that $f(\emptyset)\neq\emptyset$. Foundational ideas, including axioms or tautologies, can be generated without prior context. That case is aligned with LLMs possessing basic knowledge from pretraining that can be activated by minimal inputs. Then, with foundational knowledge given by:
$K_{\emptyset}=\bigcup_{f\in\mathcal{F}_{C}}f(\emptyset),$ (47)
we define the context-relevant knowledge as:
$\mathcal{R}_{C}(\emptyset)=K_{\emptyset}.$ (48)
When the context is empty, the only relevant knowledge is the foundational knowledge. Then, $\mu_{C}(A|\emptyset)>0$ if and only if $A\cap K_{\emptyset}\neq\emptyset$. Reasoning without context only expands the knowledge set $A$ if the conclusions $f(A)$ represent foundational knowledge. Finally, $\mathcal{E}_{C}(\emptyset|\emptyset)=K_{\emptyset}$.
We permit structures where $\mu_{C}$ is not monotonic in context. Consider, $\mu_{C}(A|Q)=|A\setminus Q|$ and $\mathcal{R}_{C}(Q)=\mathcal{K}\setminus Q$. Let us then see what happens if $A=\{k\}$, $Q_{1}=\emptyset$ and $Q_{2}=\{k\}$. We have:
$Q_{1}\subseteq Q_{2}\quad\text{and}\quad\mu_{C}(A|Q_{1})=|\{k\}\setminus\emptyset|=1>0=|\{k\}\setminus\{k\}|=\mu_{C}(A|Q_{2}).$ (49)
Increasing the context may reduce the information contributions. If the extended context $(Q_{2})$ provides an alternative explanation for phenomena related to original context $(Q_{1})$, then marginal contribution of available knowledge $(A)$ may decrease. There is a nice interpretation of that property, that we all may recognize very well. Often when we learn about some fundamental concept we overestimate our deep understanding of it and we tend to develop overconfident belief that we have become experts in a domain. But the more we learn, the less confident we feel, eventually coming to the conclusion that we know that we know nothing (or not enough at least), following the steps of Socrates.
@@ 4.2 Statements, Queries, and Ground Truth
Queries and responses are statements expressed in natural language that LLM is able to process and generate. We need to compare those statements and relate them to knowledge. It is also useful to distinguish queries from responses by their nature, respectively, interrogative or declarative. Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions).
We connect the abstract knowledge space $\mathcal{K}$ to the linguistic inputs and outputs of the model. We model the language space analogously to the knowledge space.
Definition 9 (Statement Space and Knowledge Mapping).
The space of natural language statements $(\mathcal{L},d_{\mathcal{L}})$ is a Polish space. The query space $\mathcal{Q}$ and the response space $\mathcal{R}$ are Borel subsets of $\mathcal{L}$. The knowledge mapping $K\colon\mathcal{L}\rightarrow\mathcal{B}(\mathcal{K})$ is a Borel measurable function that maps statements to the knowledge they reference.
We require the output of $K$ to be in the Borel sets $\mathcal{B}(\mathcal{K})$, because $K(q)$ serves as the context for $\mathcal{E}_{C}(A|Q)$. As established earlier (Theorem 2), the context $Q$ must be Borel to guarantee that the reasoning process remains closed within the analytic sets representing knowledge.
Before we define the ground truth mapping, we need to acknowledge its epistemological limitations. We recognize that the very idea of ground truth validation is limited (or subjective at least). It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query. Therefore, in practice that inaccessibility of the complete truth validation function (even in the idealized and rigorous mathematical reasoning) is one of the reasons why hallucination cannot be eliminated entirely (cf. kalai2025why ). High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth.
Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping.
@@
The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{B}(\mathcal{K})$ is a Borel measurable function that associates each query $q\in\mathcal{Q}$ with the Borel subset of relevant knowledge labeled as (true) correct answer.
In other words, $T(q)$ represents our best available approximation of (labeled as) true relevant knowledge. There is some external agency providing a data base or dataset telling what is and what is not true response to a query, and that evaluation has a form of a label assigned to knowledge elements. Similar to knowledge mapping $K(q)$, ground truth $T(q)$ contributes to the context $Q=K(q)\cup T(q)$, and thus must also be Borel to ensure $Q\in\mathcal{B}(\mathcal{K})$.
We focus our analysis on non-trivial queries that require the integration of diverse information sources.
@@
A query $q\in\mathcal{Q}$ is non-trivial at budget $C$ if there exist knowledge subsets $A$ and $B$ such that for the context $Q=K(q)\cup T(q)$:
@@
A query is non-trivial when at least two agents disagree on some fact that is both encoded and available, and relevant to answering that query. Non-triviality implies the existence of relevant, accessible disagreement among agents.
Example.
Consider a query $q$:
tell me about recent AI breakthroughs.
Its context is given by $Q=K(q)\cup T(q)$. Then, the relevant knowledge $\mathcal{R}_{C}(Q)$ includes facts derivable within $C$ computational steps that relate to AI advances. Non-triviality occurs when different agents (e.g., attention heads) encode conflicting information about these breakthroughs. ∎
Definition 11 describes the ability of a response generation system to store knowledge that is necessary for the impossibility to emerge. In other words, for trivial queries the impossibility results do not hold. We may view that ability to respond to non-trivial queries as an analog of the formal system complexity in the Goedel’s framework that gives rise to incompleteness of consistent formal systems godel1931formal .
4.3 Auction-Theoretic Model of Inference
There are many ways LLM can generate answers to a query. By design, it is a controlled random process governed by fine-tuned probability distributions. We are going to model that process as an auction of ideas in which agents (representing competing ideas is the space of all possible knowledge representations) use their privately encoded knowledge (internal states) to construct bids (candidate responses) and influence the response.
We first define the space representing the internal states (e.g., parameters or activations) of the bidding agents, and the mapping from these states to the knowledge they encode.
Definition 12 (Private Knowledge Configuration).
The private knowledge configuration of agent $i$, denoted $\Theta_{i}$, lives in a Polish space. The joint space is the Cartesian product $\Theta=\prod_{i=1}^{n}\Theta_{i}$. The configuration mapping $\kappa\colon\Theta\mapsto(\mathcal{A}(\mathcal{K}))^{n}$ is a measurable function, where $\kappa_{i}(\theta_{i})\in\mathcal{A}(\mathcal{K})$ is the analytic knowledge set associated with agent $i$’s configuration $\theta_{i}$ of privately encoded knowledge.
For each agent, $\theta_{i}$ is the inherent configuration of encoded knowledge that uniquely determines its capabilities and behavior, independent of the specific input context being processed. In a trained transformer model, that may correspond to the learned parameters (weight matrices) of attention head:
$\theta_{h}=\left(W_{Q}^{(h)},W_{K}^{(h)},W_{V}^{(h)},W_{O}^{(h)}\right),$ (51)
where $W_{Q}^{(h)}$ and $W_{K}^{(h)}$ are query and key weights, $W_{V}^{(h)}$ and $W_{O}^{(h)}$ are value and output weights. Private knowledge configuration $\theta_{h}$ tells what the agent is looking for and where, and how it extracts the relevant information.
@@ 4.3 Auction-Theoretic Model of Inference
  • strategy space $S=S_{1}\times\cdots\times S_{n}$, where $S_{i}$ is a Polish space (representing bids, e.g., outputs of attention heads or probability distributions),
  • a measurable outcome function $g\colon S\mapsto\mathcal{R}$ mapping strategy profiles $s\in S$ to responses in the Polish response space $\mathcal{R}$,
  • a measurable information contribution function $p\colon S\times\Theta\mapsto\mathbb{R}^{n}$ determining how strategy profile $s\in S$ inference.
A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge in response to a query. It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction. Examples of strategies include activations deciding which tokens to attend or how strongly to activate neurons for different patterns. Then, given a set of submitted bids (in the form of strategy profile) $s=(s_{1},\dots,s_{n})$, the response mechanism uses outcome function $g(s)$ to select and generate response. One straightforward example is a softmax function in LLM’s final layer. Finally, given a knowledge configuration profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, the information contribution function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output. In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
@@
The inference process involves $n$ agents, indexed by $\mathcal{N}=\{1,\dots,n\}$. Each agent chooses a strategy $s_{i}\in S_{i}$ to maximize a quasi-linear utility:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta).$ (52)
The valuation function $v_{i}\colon\mathcal{R}\times\Theta_{i}\mapsto\mathbb{R}_{+}$ is measurable, representing the agent’s preference for the outcome $g(s)$ given its private knowledge configuration $\theta_{i}$.
One way to think about the concept of agents is that it represents different attention heads in transformer models, or components of mixture-of-experts architectures, or activation paths fired by the same query. Each agent’s private knowledge represents model parameters utilized by an agent to construct a response. The utility function can be best described as a partial contribution of a candidate response to LLM loss function minimization. It is designed to explain why any inference path is activated in response to the LLM prompt.
The agents represent algorithms and architectural components of any LLM implementation. That is an analog of the Revelation Principle in game theory. If we know strategies (of using knowledge) the players (conscious beings with individual goals) will use in a game, we can implement those strategies as an algorithm (or agent) with parameters representing private preferences or knowledge of the players. In the case considered, we are the players selecting architecture and training the LLM.
The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs.
@@
Justification of the independent private knowledge model
For the idealized analysis in Theorem 6, we require that the agent’s valuation depends only on its own private knowledge configuration and that the configurations are statistically independent. That is restrictive, but it allows us to establish a baseline impossibility, which is then generalized in subsequent proofs that accommodate correlated beliefs.
In the realistic settings of transformers, heads may share layer-norm signals, residual streams, and optimizer state. That may result in correlated (or, so called, affiliated) inference signals between token representations. Furthermore, the superposition hypothesis, introduced in mechanistic interpretability studies, suggests cooperative error-canceling due to strong correlations.
We show that impossibility arises in both independent and correlated settings. We refer to Green-Laffont theorem green1977characterization in the independent private knowledge setting. Then, we refer to the truthful probability elicitation and proper scoring theory introduced by Savage savage1971elicitation , which does not require independent private knowledge.